# Vulnerable Connectivity Caused by Local Communities in Spatial Networks

### v2

@@ 3.2 Robustness of Connectivity Against Recalculated Betweenness Attacks
Robustness of connectivity plays a crucial role in complex networks for quantifying a network’s ability to maintain the functionality when subjected to intentional attacks or random failures, such as RB, ID and RF mentioned in subsection 2.2. As the measures, we use the robustness index $R$ [34] and the critical fraction $q_{c}$ [35]. Here, the robustness index $R$ is the area under the curve of the LCC’s relative size $S^{1st}(q)/N$ against the fraction $q$ of removed nodes. A higher $R$ value indicates greater overall robustness of connectivity. The critical fraction $q_{c}$ is the value of $q$ at which the relative size $S^{2nd}(q)/N$ of the second LCC reaches its peak. It captures the critical point of structural change in a network. A higher $q_{c}$ value suggests that the network can withstand more removals of nodes before substantial fragmentation. However, $R$ is a more precise measure because different $R$ values are possible for rapidly and gradually decreasing curves of $S^{1st}(q)/N$ even with a same $q_{c}$. The combination of these two measures provides us with ways to observe not only how tenaciously a network maintains its connectivity against attacks but also when it reaches its breaking point.
Figures 7 show the remarkable relation between modularity $Q$ and the robustness measures $R$ against RB attacks. As $Q$ increases (with stronger community structure), $R$ decreases (with weaker robustness of connectivity) for all seven major Japanese areas with different node’s locations. S36 Figure and S37 Figure in Supporting Information provide additional evidences of similar results in the networks with sizes $N=100$ and $N=10000$. In other words, as higher $Q$ values with stronger community structure, the robustness becomes more vulnerable in the spatial network because of the limited number of inter-links (especially against RB attacks whose targets are the end-nodes of such bridges between communities).
@@
We further investigate how community structure weakens the robustness of connectivity for these networks in explaining the details for examples of Tokyo 1024 node’s networks. The results show that the spatial concentration of nodes creates strong community structure, which weaken the robustness. In Figures 8 and 9, it is common that both randomized networks and 2DL networks show rightward shifts in their curves as improving the robustness to the original networks. When $S^{2nd}(q)/N$ reaches its peak at $q_{c}$, the corresponding $S^{1st}(q)/N$ curve is rapidly decreased. We explain the details of comparisons in the following, where the $R$ value is represented by the area under $S^{1st}(q)/N$ curve, and that the $q_{c}$ value is indicated by the peak of $S^{2nd}(q)/N$ curve.
@@
Figure 8 shows the robustness curves comparing the original networks (Pop., Inv., and Uni. node’s connected by RNG or GG) with their degree-preserved randomized (rewired) counterparts where communities are eliminated. By comparing RNG (Figures 8ac) and GG (Figures 8bd), we confirm that all curves in GG exhibit rightward shifts compared to their corresponding curves for RNG with same colors and line types. This rightward shift indicates higher robustness for GG as shown by both larger areas under the $S^{1st}(q)/N$ curves (higher $R$ values) and delayed peaks in $S^{2nd}(q)/N$ curves (higher $q_{c}$ values). Remember that RNG has fewer links than GG as the subgraph.
@@
To investigate the effect of community structure on the robustness of connectivity, we compare the original networks (represented by RGB solid lines) with their randomized counterparts (shown as RGB dashed lines) in Figure 8a-d, where networks denoted by a same color line have an identical degree distribution $P(k)$ as shown in Figure 10. Green and red dashed lines (randomized Inv.) show significant rightward shifts compared to the corresponding solid lines (the original Pop., Inv.). These lines indicate that weaker community structure enhances the robustness. Similarly, blue dashed lines (randomized Uni.) show the improvement of robustness to its solid counterparts. In each of Figure 8a-d for the original networks (RGB solid lines), both green (Pop.) and red (Inv.) solid lines exhibit leftward shifts compared to blue lines (Uni.). It indicates that Uni. networks have higher robustness. These comparisons consistently show that stronger community structure (higher modularity) lead to lower robustness, also as evidenced by increasing modularity Q values (see the differences for the original and randomized networks in Table 1). In other words, the limited number of inter-links create potential vulnerabilities against attacks. However, this relation becomes slightly complex in randomized networks with some exceptions: For RNG (Figure 8a), the green dashed lines (randomized Pop.) exhibit higher robustness than the blue dashed lines (randomized Uni.), while the red dashed lines (randomized Inv.) show lower robustness. For GG (Figure 8b), both green and red dashed lines have lower robustness than the blue dashed lines. With the exception of randomized RNG networks, these cases suggest that a uniform distribution of nodes contributes to improved the robustness in Figure 8ab.
@@
Figure 9 shows the robustness curves comparing the original networks with weaker community networks where nodes are relocated onto 2D lattices under same degree distributions of Pop., Inv., and Uni., respectively, to investigate how spatial concentrations of nodes affect the robustness of connectivity. In Figure 9a-d, the cyan solid lines (2DL-Uni.) show higher robustness than the blue solid lines (the original Uni.) as evidenced by both larger area under the $S^{1st}(q)/N$ curves (higher $R$ value) and delayed peak in $S^{2nd}(q)/N$ curve (higher $q_{c}$ value). Similar improvements of robustness are also observed, when comparing the yellow solid lines (2DL-Pop.) with the green solid lines (the original Pop.), and comparing the magenta solid lines (2DL-Inv.) with the red solid lines (the original Inv.). When comparing these networks, we observe a clear hierarchy in the robustness that correlates with the strength of community structure. As shown in Table 1, the original networks show stronger community structure while 2DL networks show weaker community structure. As exceptions in both randomized and relocated networks, RNG-Pop networks (in bold) have lower modularity $Q$ than RNG-Uni networks, which may be attributed to RNG-Pop networks have more concentrated degree distributions than RNG-Uni networks. In Figures 8 and 9, randomized networks (RGB dashed lines) without communities have the highest robustness, followed by 2DL networks (Cyan-Yellow-Magenta, CYM solid lines), while the original networks (RGB solid lines) with strong communities have the lowest robustness. For instance, the blue dashed lines (randomized Uni.) show to be higher robust than the cyan solid lines (2DL-Uni.), which show to be even more robust than the blue solid lines (the original Uni.). Similar results can be observed in the comparisons between solid-green and dashed-yellow lines (Pop.), and between solid-red and dashed-magenta lines (Inv.). These comparisons reveal that weaker community structure through more homogeneous node’s locations improves the robustness as shown by the rightward shifts of CYM solid lines compared to that in the original networks. In summary, spatial networks with weaker community structure have higher robustness against RB attacks.
@@ 3.3 Robustness of Connectivity Against Initial Degree Attacks and Random Failures
We further investigate the robustness of connectivity against typical Initial Degree (ID) attacks and Random Failures (RF). For Uni. networks, we confirm that $q_{c}$ values are almost identical to the analytical values [36] as shown in Tables 2 and 3. Table 2 shows the robustness index $R$ and critical fraction $q_{c}$ against ID attacks, while Table 3 shows the corresponding $R$ and $q_{c}$ values against RF. Note that a clear hierarchy exists as following the relation $R^{RB}<R^{ID}<R^{RF}$ and ${q_{c}}^{RB}<{q_{c}}^{ID}<{q_{c}}^{RF}$ (see S2 Table in Supporting Information for detailed comparison) in the robustness across different attack strategies.
In both ID attacks and RF, networks whose nodes are located according to Pop. and Inv. show lower $R$ and $q_{c}$ values than those to Uni. networks as similar to the results against RB attacks except some cases. As values with upper-triangles ($\triangle$) in Tables 2 and 3, some networks based on Pop. and Inv. have higher robustness than those on Uni.. The reason may be caused by two factors: (1) Higher average degree $\langle k\rangle$ in RNG Pop. networks as shown in Table 4, which may provide additional alternative paths after node removals; (2) The presence of regular grid-like parts in certain regions of Pop. networks contributes to enhance the robustness against RF as shown in S15 Figure. The effect of grid-like parts is explained as follows. In our study, all nodes are located limitedly at the centers of mesh blocks (500m × 500m), rather than having arbitrary positions in a geographical space freely. When nodes are densely distributed in neighboring mesh blocks, as often occurs in Pop. networks, they tend to form grid-like parts. We quantify these grid-like parts by calculating the ratio of nodes that have degree 4 with four neighbors of degree 4, which represents a typical local grid formation. As shown in Table 5, Pop. networks show notably higher ratio of such grid-like parts (ranging from $2.05\%$ to $15.14\%$ in RNG and $2.25\%$ to $15.33\%$ in GG) compared to Inv. networks (mostly $0\%$ in RNG and below $1.17\%$ in GG) and Uni. networks ($0\%$ in RNG and $0.79\%$ in GG) as values with upper-triangles ($\triangle$). According to percolation theory [37], a 2D square lattice network maintains its connectivity until the fraction $0.4073$ (percolation threshold $pc=0.5927$) of randomly removed nodes. This critical fraction $0.4073$ is larger than the analytical values $0.205$ in RNG and $0.365$ in GG shown in Table 3. The presence of these grid-like parts increases the critical fraction than the above values in RNG and GG on free node’s locations at random.
4 Conclusion
Our study investigates the impact of local community structure on the vulnerability of spatial networks with concentration of nodes. The node’s locations are according to the order of real population (Pop.), inverse real population (Inv.), and uniformly at random (Uni.), while the links between nodes are short by using proximity graphs of RNG and GG as modeling road and communication systems, respectively. Remember that Pop., Inv., and Uni. are considered to investigate the effect of strengths of community structure on the robustness. We observed the emergence of local communities in Pop. and Inv. networks connected by short links (see Figs. 3 and 4), which weakens the networks by removing small number of inter-community bridge links (see Figs. 5 and 6). In the details, for the measures of robustness index $R$ and critical fraction $q_{c}$, Pop. and Inv. networks are weaker than Uni. networks against intentional attacks and random failures in both RNG and GG, because Pop. and Inv. networks have stronger local communities than Uni. networks (see Figs. 8 and 9). In particular, both $R$ and $q_{c}$ decrease with increasing modularity $Q$ (see S13 Figure, S14 Figure and Fig. 7). Therefore, we conclude that the emergence of local communities weakens the robustness of connectivity in spatial networks.
For this negative impact of local communities on the robustness, it is important to understand how to reduce their effects. Local communities emerge from the spatial concentration of nodes connected by short links. While the strength of local communities could be mitigated by either distributing nodes more evenly or creating long links, the former approach may be impractical due to high land acquisition costs. Therefore, strategically establishing long links offers a more feasible solution to balance reliability and construction costs.
However, our study has several limitations. For example, while we focus on road and communication networks, the applicability of our findings remains to other infrastructure networks, such as power grids and water supply systems. Moreover, although we thoroughly investigated the effects of RB, ID, and RF attacks, the results of other more distractive attacks, such as belief propagation attacks [38] and collective influence attacks [39], on the vulnerability of spatial networks warrants further investigation. In addition, the universal applicability of our conclusions to regions outside Japan requires additional verification through studies in other countries and regions.
List of Abbreviations
RGB
Red-Green-Blue (color scheme)
CYM
Cyan-Yellow-Magenta (color scheme)
GG
Gabriel Graph
RNG
Relative Neighborhood Graph
2DL
Two-Dimensional Square Lattice
Pop.
Population
Inv.
Inverse Population
Uni.
Uniform
RB
Recalculated Betweenness (attack strategy)
ID
Initial Degree (attack strategy)
RF
Random Failures
SF
Scale-Free
ER
Erdős-Rényi
SPP
Spatial Point Process
LCC
Largest Connected Component
Figure 1: Illustration of connection constraints for (a) RNG and (b) GG. A link colored by red is established between two nodes colored by blue, when no other node exists within the shaded area.
Figure 2: Coverage diagram of radio waves between two base stations in wireless communication. The ranges of strong beams are shown by blue and orange shades. The center circle represents the signal interference area. If other base stations exist within it, the two stations cannot be connected.
Figure 3: Visualization of community structure in Tokyo before node removal. $N=1024$ nodes are located by the decreasing order of population (Pop.). Different colors represent different communities estimated by Louvain method. There are clear community formations particularly in densely populated areas.
Figure 4: Visualization of community structure in Tokyo before node removal. $N=1024$ nodes are located by the inverse order of population (Inv.). Different colors represent different communities estimated by Louvain method. There are different community formations compared to Figure 3.
Figure 5: Tokyo networks with $N=1024$ nodes after 10% node removals by recalculated betweenness (RB) attacks. The locations of nodes are selected by decreasing order of population (Pop.). The fragmentation is observed by removing links between communities especially in densely populated areas
Figure 6: Tokyo networks with $N=1024$ nodes after 10% node removals by recalculated betweenness (RB) attacks. The locations of nodes are selected by inverse order of population (Inv.). The fragmentation is observed by removing links between communities which are different as compared to Figure 5.
Figure 7: Relation between robustness index $R^{RB}$ and modularity $Q$ in networks with $N$ = 1024 nodes
Figure 8: Robustness against recalculated betweenness (RB) attacks for Tokyo networks with $N=1024$ nodes. For rewired (Randomized networks) lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) the relative size $S^{1st}(q)/N$ of largest connected component, and (c) (d) the critical fraction $q_{c}$ at the peak of the relative size $S^{2nd}(q)/N$ of second largest component.
Figure 9: Robustness against recalculated betweenness (RB) attacks for Tokyo networks with $N=1024$ nodes. For 2DL (relocated networks) lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) the relative size $S^{1st}(q)/N$ of largest connected component, and (c) (d) the critical fraction $q_{c}$ at the peak of the relative size $S^{2nd}(q)/N$ of second largest component.
Figure 10: Similar degree distributions of bell-shapes for Pop., Inv., and Uni. in Tokyo networks with $N$ = 1024 nodes.
Cities Original Randomized Relocated
RNG GG RNG GG RNG GG
Inv. Pop. Inv. Pop. Inv. Pop. Inv. Pop. Inv. Pop. Inv. Pop.
Fukuoka 0.8905 0.8953 0.8728 0.8857 0.786 0.6978 0.5776 0.661 0.8726 0.8298 0.7677 0.8269
Hiroshima 0.8866 0.8993 0.8717 0.8898 0.7863 0.7001 0.5815 0.6771 0.8704 0.8332 0.7662 0.8203
Keihan 0.8903 0.8924 0.8651 0.8777 0.7874 0.6888 0.5861 0.6383 0.8693 0.8302 0.7756 0.8149
Nagoya 0.8880 0.8929 0.8612 0.8764 0.7969 0.6927 0.595 0.6412 0.8807 0.8408 0.7744 0.8158
Tokyo 0.8846 0.8814 0.8615 0.8610 0.7895 0.682 0.5741 0.6362 0.8644 0.8301 0.7548 0.8177
Sendai 0.8853 0.8908 0.8612 0.8801 0.7824 0.6969 0.58 0.6611 0.8653 0.8387 0.7648 0.8184
Sapporo 0.9017 0.8775 0.8875 0.8722 0.8019 0.6528 0.6118 0.6335 0.8745 0.8187 0.7863 0.8144
Uniform 0.8735 0.8557 0.7704 0.5451 0.8649 0.745
Table 1: Modularity $Q$ in networks with 1024 nodes for seven major Japanese areas under three different conditions: original networks, after randomization (rewired), and after nodes relocation on a 2DL (relocated). Higher values indicate stronger community structure. Values with upper-triangles ($\triangle$) indicate where the cases of Pop. and Inv. have higher modularity than the case of Uni., while values in bold indicate lower modularity than the case of Uni. for both RNG and GG.
@@ 4 Conclusion
Table 2: Robustness index ($R^{ID}$) and Critical fraction ($q_{c}^{ID}$) against Initial Degree (ID) attacks in the original networks with $N=1024$ nodes for seven major Japanese areas. Higher values indicate greater robustness. Values with upper-triangles ($\triangle$) indicate where the cases of Pop. and Inv. have stronger robustness than the case of the Uni. network.

### v4

@@ 3.2 Robustness of Connectivity Against Recalculated Betweenness Attacks
Robustness of connectivity plays a crucial role in complex networks for quantifying a network’s ability to maintain the functionality, when subjected to intentional attacks or random failures such as RB, ID and RF mentioned in subsection 2.2. As the measures, we use the robustness index $R$ [61] and the critical fraction $q_{c}$ [62]. Here, the robustness index $R$ is the area under the line of the LCC’s relative size $S^{1st}(q)/N$ against the fraction $q$ of removed nodes. A higher $R$ value indicates greater overall robustness of connectivity. The critical fraction $q_{c}$ is the value of $q$ at which the relative size $S^{2nd}(q)/N$ of the second LCC reaches its peak. It captures the critical point of structural change in a network. A higher $q_{c}$ value suggests that the network can withstand more removals of nodes before substantial fragmentation. However, $R$ is a more precise measure because different $R$ values are possible for rapidly and gradually decreasing lines of $S^{1st}(q)/N$ even with a same $q_{c}$. The combination of these two measures provides us with ways to observe not only how tenaciously a network maintains its connectivity against attacks but also when it reaches its breaking point.
To understand the effect of community structure on robustness of connectivity, we compare the original networks (triangles) with the 2D lattice networks (crosses), which relocate the original networks’ nodes on a 2D lattice grid by removing spatial concentrations of nodes but maintaining the degree distributions for node’s locations based on Pop. (green and yellow), Inv. (red and magenta), and Uni. (blue and cyan) in 7 Japanese areas. Fig. 5 shows clear monotone decreasing between modularity $Q$ and robustness index $R^{RB}$ against RB attacks in these networks. In both GG and RNG, the original Pop.-based and Inv.-based networks show higher $Q$ and lower $R^{RB}$ than those in Uni.-based networks by comparing triangle marks with three colors (green, red, and blue). Such results are caused by that the difference of Pop., Inv., and Uni. is the spatial concentration of nodes under the similar degree distributions $P(k)$ as shown in Fig. 6. Moreover, by comparing cross and triangle marks for each of the node’s locations based on Pop., Inv., and Uni., respectively, in the original and 2D lattice networks under the same degree distributions, we show that 2D lattice networks (cross marks with yellow, magenta, and cyan) have lower $Q$ and higher $R^{RB}$ than the original networks (triangle marks with green, red, and blue) because community in 2D lattice networks have been weakened by relocating nodes on a 2D lattice with longer rewiring links, thus improve the robustness index $R^{RB}$. We note that small variations for 7 areas in each of the original and 2D lattice networks with the node’s locations based on Pop. and Inv. because of slightly different $P(k)$ as shown in Fig. 6. However, an exception appears in RNG, where 2D lattice Pop.-based $<$ Uni.-based $<$ Inv.-based networks for $Q$, and the relation for $R^{RB}$ has the inverse order. This exception may caused by the slight different degree distribution $P(k)$, as shown in Fig. 6. In summary, as higher $Q$ values with stronger local community, the robustness becomes more vulnerable in the spatial network because of the limited number of inter-links (especially against RB attacks whose targets are the end-nodes of such bridges between communities). Supporting Informations S36 Fig and S37 Fig confirm these relations for networks with $N=100$ and $N=10000$.
Fig. 5 Modularity $Q$ and Robustness Index $R^{RB}$. Relation between modularity $Q$ and robustness index $R^{RB}$ in networks with $N=1024$ nodes of 7 Japanese areas against RB attacks. Each point represents the values of $R^{RB}$ and $Q$ for the original (triangles) and 2D lattice (crosses) networks with node’s locations based on Pop. (green/yellow), Inv. (red/magenta), and Uni. (blue/cyan). Links are created by RNG (a) or GG (b). Inset in (a) enlarges the lower-right area with densely points. A monotone decreasing is observed: networks with stronger communities (higher $Q$) tend to have lower robustness index $R^{RB}$. Similar relations are observed in RNG, while the ranges of $Q$ and $R^{RB}$ differ slightly.
Fig. 6 Degree Distributions. Degree distributions $P(k)$ in the original Tokyo networks with $N=1024$ and with node’s locations based on Pop., Inv., and Uni.. All cases have bell-shaped forms with peaks around $k=2$ to $4$.
For the reason that the robustness of connectivity becomes more vulnerable as stronger communities, we investigate a relation between robustness of connectivity and the following sparsity index $SI(G_{w})$ [63] [64]. The $SI(G_{w})$ is defined as follows:
$SI(G_{w})=1-\frac{1}{N^{2}T_{1}}\sum_{j=1}^{k}w_{j}f_{j}\left(f_{j}+2f_{j+1}+2f_{j+2}+\dots+2f_{k}\right)$
Here, the total link weight $T_{1}=\sum_{e=1}^{M^{\prime}}w_{e}f_{e}$, link weight $w_{e}$ is defined by normalized Euclidean distances between nodes connected by a link $e$, and $f_{e}$ is its frequency. The parameter $M^{\prime}$ denotes the number of different link weights observed in the network. The normalization factor $N^{2}T_{1}$ ensures that $SI(G_{w})$ is bounded within the range $[0,1]$. Fig. 7 shows that $R^{RB}$ in both RNG and GG decreases monotonically with respect to $SI(G_{w})$. In other words, networks with higher spatial sparsity index $SI(G_{w})$ tend to have lower robustness index $R^{RB}$. The Pearson’s correlation coefficients $r<0$ for all cases against RB attacks confirmed these monotone decreasing as shown in supporting information S13 Table. The blue, red, and green points correspond to Uni.-based, Inv.-based, and Pop.-based networks in 7 Japanese areas. Note that the robustness index $R^{RB}$ on the vertical axis varies within a narrow range, while the sparsity index $SI(G_{w})$ on the horizontal axis spans a much wider range. Because in Pop.-based networks, most nodes are densely concentrated in urban regions, but a few nodes are located in peripheral regions. These distant nodes connect to the urban regions by long links, which significantly increase the total link weight $T_{1}$. Thus, the sparsity index $SI(G_{w})$ becomes larger. In Inv.-based networks, nodes are positioned more evenly than in Pop.-based networks (see Fig. 3ab and 4ab for comparison), which results in fewer long links therefore a lower $T_{1}$ and $SI(G_{w})$. In Uni.-based networks, node’s locations are the most uniformly positioned compared to Pop.-based and Inv.-based networks, yielding the lowest $T_{1}$ and the lowest $SI(G_{w})$.
Fig. 7 Robustness Index $R^{RB}$ and Sparsity Index $SI(G_{w})$. Relation between robustness index $R^{RB}$ and sparsity index $SI(G_{w})$ for networks with $N=1024$ nodes in seven Japanese areas. Each point represents the results for (a) RNG and (b) GG with node’s locations based on Pop. (green), Inv. (red), and Uni. (blue). Colored points show clear monotone decreasing that networks with higher sparsity index $SI(G_{w})$ tend to have lower robustness index $R^{RB}$. We note that the robustness index $R^{RB}$ on the vertical axis varies within a narrow range, while the sparsity index $SI(G_{w})$ on the horizontal axis spans a much wider range.
In Fig. 5, we observe that higher modularity $Q$ is associated with lower robustness index $R$. Similarly, in Fig. 7, we note that networks with higher spatial sparsity index $SI(G_{w})$ tend to have lower robustness index $R$. These relations suggest a possible relation between sparsity index $SI(G_{w})$ and modularity $Q$. Thus, we investigate the relation between $SI(G_{w})$ and $Q$ as shown in Fig. 8 with green, red and blue points corresponding to Pop.-based, Inv.-based, and Uni.-based networks with $N=1024$ nodes in seven Japanese areas. There is clear monotone increasing between $SI(G_{w})$ and $Q$ of these colored points. In other words, higher sparsity index $SI(G_{w})$ is associated with higher modularity $Q$. The Pearson’s correlation tests confirm the relation in both RNG and GG with the significance (see Supporting Information S14 Table for the detail). We note that the scales of horizontal ($SI(G_{w})$) and vertical ($Q$) axes are quite different. There are a large range of $SI(G_{w})$ and a small range of $Q$. In particular, it is remarkable that Pop.-based $>$ Inv.-based $>$ Uni.-based networks for $SI(G_{w})$, while the values of $Q$ are nearly same as high. This ranking of $SI(G_{w})$ reflects the differences in node’s locations, as previously discussed in the analysis of Fig. 7.
Fig. 8 Sparsity Index ($SI(G_{w})$) and Modularity $Q$. Relation between sparsity index $SI(G_{w})$ and modularity $Q$ for the networks with $N=1024$ nodes in seven Japanese areas. Each point represents the result for (a) RNG and (b) GG with node’s locations based on Pop. (green triangles), Inv. (red squares), or Uni. (blue circle). Colored points show clear monotone increasing: networks with higher spatial sparsity index $SI(G_{w})$ tend to have higher $Q$. Pop.-based networks (green triangles) have the highest $SI$ values because of long-range link between spatially concentrated urban regions and peripheral regions, while Uni.-based (blue circle) has the lowest $SI$ because of uniformly positioned node’s locations. We note that the scales of vertical and horizontal axes are quite different. In other words, the vertical differences are ignorable.
@@
We further investigate how community structure weakens the robustness of connectivity for these networks in explaining the details for examples of Tokyo networks with $N=1024$ nodes. The results show that the spatial concentration of nodes creates strong community structure, which weaken the robustness. In Figs. 9 and 10, it is common that both randomized networks and 2DL networks show rightward shifts in their lines as improving the robustness to the original networks. When $S^{2nd}(q)/N$ reaches its peak at $q_{c}$, the corresponding $S^{1st}(q)/N$ line is rapidly decreased. We explain the details of comparisons in the following, where the $R$ value is represented by the area under $S^{1st}(q)/N$ line, and that the $q_{c}$ value is indicated by the peak of $S^{2nd}(q)/N$ line.
Fig. 9 Robustness Against RB for the Original and Randomized Tokyo Networks. Robustness against recalculated betweenness (RB) attacks in Tokyo networks with $N=1024$ nodes. For rewired (randomized networks) lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) the relative size $S^{1st}(q)/N$ of largest connected component, and (c) (d) the critical fraction $q_{c}$ at the peak of the relative size $S^{2nd}(q)/N$ of second largest component.
Fig. 10 Robustness Against RB for the Original and Relocated Tokyo Networks. Robustness against recalculated betweenness (RB) attacks in Tokyo networks with $N=1024$ nodes. For 2DL (relocated networks) lines, the rewiring process preserves the original degree distributions. Two measures are applied: (a) (b) the relative size $S^{1st}(q)/N$ of largest connected component, and (c) (d) the critical fraction $q_{c}$ at the peak of the relative size $S^{2nd}(q)/N$ of second largest component.
@@
Fig. 9 shows the robustness in comparing the original networks with node’s locations based on Pop., Inv., and Uni. to their randomized counterparts under the same degree distributions $P(k)$. By comparing RNG (Figs. 9ac) and GG (Figs. 9bd), we confirm that all lines in GG exhibit rightward shifts compared to their corresponding lines for RNG with same colors and line types. This rightward shift indicates higher robustness for GG as shown by both larger areas under the lines of $S^{1st}(q)/N$ (higher $R$ values) and delayed peaks in lines of $S^{2nd}(q)/N$ (higher $q_{c}$ values).
@@
To investigate the effect of community structure on the robustness of connectivity, we compare the original networks (represented by RGB solid lines) with their randomized counterparts (shown as RGB dashed lines) in Fig. 9a-d. As shown in Fig. 6, the original networks with three node’s locations based on Pop., Inv., and Uni. have similar bell-shaped degree distributions with a peak around $k=2$ to $k=4$. These bell-shaped degree distributions contrast to Scale-Free networks with high-degree hubs by preferentially minimizing the number of hops between nodes in the networks [65]. However, in our study, the networks are constructed using proximity graphs such as RNG and GG, where nodes are connected to its relatively neighbors without long-range shortcuts. In planar graphs, node degrees are generally bounded and tend not to exceed six [37, page 10]. Therefore node degrees are constrained in a narrowly range, and the absence of high degree hubs produce a bell-shape degree distribution in contrast to the heavy-tailed degree distributions in Scale-Free networks. Green and red dashed lines (randomized Inv.-based networks) in Fig. 9 show significant rightward shifts compared to the corresponding solid lines (the original Pop.-based, Inv.-based networks). These lines indicate that weaker community structure enhances the robustness. Similarly, blue dashed lines (randomized Uni.-based networks) show the improvement of robustness to its solid counterparts. In each of Fig. 9a-d for the original networks (RGB solid lines), both green (Pop.-based networks) and red (Inv.-based networks) solid lines exhibit leftward shifts compared to blue lines (Uni.-based networks). It indicates that Uni.-based networks have higher robustness. These comparisons consistently show that stronger community structure (higher modularity) lead to lower robustness, also as evidenced by increasing modularity $Q$ values (see the differences for the original and randomized networks in Table 1). In other words, the limited number of inter-links cause the potential vulnerabilities against attacks. However, this relation becomes slightly complex in randomized networks with some exceptions: For RNG (Fig. 9a), the green dashed lines (randomized Pop.-based networks) exhibit higher robustness than the blue dashed lines (randomized Uni.-based networks), while the red dashed lines (randomized Inv.-based networks) show lower robustness. For GG (Fig. 9b), both green and red dashed lines have lower robustness than the blue dashed lines. With the exception of randomized networks in RNG, these cases suggest that a uniform positioning of nodes contributes to improved the robustness in Fig. 9ab.
Fig. 6 Degree Distributions. Degree distributions $P(k)$ in the original Tokyo networks with $N=1024$ and with node’s locations based on Pop., Inv., and Uni.. All cases have bell-shaped forms with peaks around $k=2$ to $4$.
@@
Fig. 10 shows the robustness in comparing the original networks with node’s locations based on Pop., Inv., and Uni. to their relocated counterparts under the same degree distributions $P(k)$, where nodes in these networks are relocated on a 2D lattice. These networks have three types of node’s locations, Pop., Inv., and Uni., to investigate how spatial concentrations of nodes affect the robustness of connectivity. In Fig. 10a-d, the cyan solid lines (2DL Uni.-based networks or 2DL-Uni.) show higher robustness than the blue solid lines (the original Uni.-based networks or Ori-Uni.) as evidenced by both larger area under the lines of $S^{1st}(q)/N$ (higher $R$ value) and delayed peak in the lines of $S^{2nd}(q)/N$ (higher $q_{c}$ value). Similar improvements of robustness are also observed, when comparing the yellow solid lines (2DL-Pop.) with the green solid lines (Ori-Pop.), and comparing the magenta solid lines (2DL-Inv.) with the red solid lines (Ori-Inv.). When comparing these networks, we observe a clear hierarchy in the robustness that correlates with the strength of community structure. As shown in Table 1, the original networks show stronger community structure, while 2DL networks show weaker community structure. As exceptions in both randomized and relocated networks, Pop.-based in RNG (in bold) have lower modularity $Q$ than Uni.-based in RNG, which may be attributed to Pop.-based networks have more concentrated degree distributions than Uni.-based networks. In Figs. 9 and 10, randomized networks (RGB dashed lines) without communities have the highest robustness, followed by 2DL networks (Cyan-Yellow-Magenta, CYM solid lines), while the original networks (RGB solid lines) with strong communities have the lowest robustness. For instance, the blue dashed lines (Randomized Uni.) show to be higher robust than the cyan solid lines (2DL-Uni.), which show to be even more robust than the blue solid lines (Ori-Uni.). Similar results can be observed in the comparisons between solid-green and dashed-yellow lines (Pop.-based networks), and between solid-red and dashed-magenta lines (Inv.-based networks). These comparisons reveal that weaker community structure through more homogeneous node’s locations improves the robustness as shown by the rightward shifts of CYM solid lines compared to that in the original networks. In summary, spatial networks with weaker community structure have higher robustness against RB attacks.
We also evaluate the significance of differences in the robustness index $R^{RB}$ and the critical fractions ${q_{c}}^{RB}$ for the original RNG and GG with three node’s locations based on Pop., Inv., and Uni. by using IBM SPSS [66] statistical tools’ one-way ANOVA tests [67] [68]. The results are shown in Supporting Information S8 Table which confirm the statistical significance of Uni.-based networks consistently for the superior robustness of connectivity than other networks, because of lower modularity $Q$ and lower sparsity index $SI(G_{w})$ as shown in Fig. 8.
@@ 3.3 Robustness of Connectivity Against Initial Degree Attacks and Random Failures
We further investigate the robustness of connectivity against typical Initial Degree (ID) attacks and Random Failures (RF). For Uni.-based networks, we confirm that $q_{c}$ values are almost identical to the analytical values [69] as shown in Tables 2 and 3. Table 2 shows the robustness index $R$ and critical fraction $q_{c}$ against ID attacks, while Table 3 shows the corresponding $R$ and $q_{c}$ values against RF. Note that a clear hierarchy exists as following the relation $R^{RB}<R^{ID}<R^{RF}$ and ${q_{c}}^{RB}<{q_{c}}^{ID}<{q_{c}}^{RF}$ (see Supporting Information S2 Table for detailed comparison) in the robustness against different attack strategies.
Table 2: $R^{ID}$ and $q_{c}^{ID}$ against $ID$ attacks in the original networks
@@
Robustness index ($R^{ID}$) and Critical fraction ($q_{c}^{ID}$) against Initial Degree (ID) attacks in the original networks with $N=1024$ nodes for seven major Japanese areas. Higher values indicate greater robustness. Values with upper-triangles ($\triangle$) indicate where the cases of Pop.-based and Inv.-based have stronger robustness than the case of the Uni.-based network.
Table 3: $R^{RF}$ and $q_{c}^{RF}$ against random failure in the original networks